# Rethinking Knowledge Distillation: A Data Dependent Regulariser With a Negative Asymmetric Payoff

### v1

@@ 3.2 Hypothesis Testing
To evaluate whether KD facilitates functional knowledge transfer, we test whether student models trained via KD are functionally more similar to the teacher than control models. Our primary hypothesis is:
KD students, on average, are no more similar to the teacher than control models.
@@
We test each functional similarity metric using a two-sided Mann-Whitney U test (significance level = 0.05). Comparisons are made between two control conditions and the variable of interest:
Same Initialisation Different Data Order (SIDDO): models with the same initialisation and architecture as the teacher, trained with seeds 10-19.
Random Control Distillation (RCD): Students trained with uniform-noise “teacher” logits (seeds 10-19; alphas 0.1, 0.5 and 0.9) (Figure 2(a)).
Standard KD (variable of interest): Students trained with real teacher logits from , using alpha values and seeds 10–19 (Figure 2(a)).
For each teacher seed, we report the mean and SEM across 10 models per condition.
@@ 4 Results and Discussion
We first examine functional transfer in small-scale settings and show that when transfer is non-marginal it is consistently asymmetric toward the teacher’s errors. We then validate these findings at larger scale on TinyImageNet, where increasing teacher train loss (via augmentation) amplifies both functional transfer and its negative asymmetry. We then demonstrate generality in negative asymmetric transfer of KD across modalities (audio and language in addition to image), show how KD can facilitate adversarial attacks and finally we provide distillation scaling experiment, in line with Busbridge et al. (2025), to show how negative asymmetric transfer is present regardless of student capacity.
Full supplemental results (datasets, architectures, and all teacher seeds) appear in the appendix: CIFAR-10 (ResNet-18, VGG19, ViT; Appendix F.2), SVHN (VGG19, ViT; Appendix F.3), audio (UrbanSound8K, SpeechCommands; Appendix G), language (Tiny Shakespeare; Appendix H), adversarial transfer (Appendix H.2), standard KD to smaller students (Appendix E), and different KD variants (Appendix C). We also show in Appendix Section B that our analysis holds for information theoretic and geometric measures alongside our functional similarity measures and that our RCD control is equivalent to label smoothing in Appendix Section D. Training details for all settings are also provided in the appendix. Unless specified otherwise, we report means and 1 SEM over 10 runs per teacher seed and condition.
4.1 Function Transfer in Small-Scale Settings (SVHN)
We begin with SVHN and ResNet18. KD yields statistically significant functional similarity at high values, but the magnitude and asymmetry of transfer vary across teacher seeds. When transfer is non-marginal, we observe a systematic increase in student–teacher agreement on incorrect predictions relative to correct ones.
Table 2 shows teacher variability: train losses of , , and with a generalisation gaps of for seeds 0, 1, and 2 respectively. Notably, the best test loss and accuracy (Table 3) are achieved by random control distillation, reducing confidence that KD’s performance gains arise from meaningful knowledge transfer and instead supporting the view of KD as a data-dependent regulariser.
@@ 4.1 Function Transfer in Small-Scale Settings (SVHN)
For the highest-train-loss teacher (seed 2), KD produces significant functional transfer across metrics at most values (Appendix Table 64; reproduced summary in Table 4), with the exception of Prediction Disagreement at . This transfer coincides with a large asymmetric payoff in prediction agreement toward the teacher’s incorrect predictions (Figure 5(a)). The lowest-train-loss teacher (seed 1) shows no significant transfer at and only partial transfer at (again, excluding Prediction Disagreement). Seed 0 (intermediate train loss) shows significant transfer at and , accompanied by asymmetric incorrect agreement (Figure 5(a)).
@@
| KD 0.5 | ✗✗✓ | ✓✗✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✓ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✗✗✓ | ✓✓✓ | ✗✗✗ | ✗✗✓ |
Across seeds, higher teacher train loss is associated with stronger (and more asymmetric) functional transfer. This is consistent with a teacher that deviates more from ground-truth labels, thereby exposing students to incorrect structure that is preferentially transferred under KD.
4.2 Function Transfer in Larger-Scale Settings
We next study TinyImageNet with ResNet50. In the base setting, KD produces significant but marginal functional gains relative to SIDDO; the corresponding prediction agreement shows no clear preference toward correct or incorrect agreement. Motivated by the SVHN analysis, we increase the teacher train loss via data augmentation (same training pipeline) – RandAugment (Cubuk et al., 2020) with the default settings – and examine the consequences for functional transfer and asymmetry.
@@ 4.2 Function Transfer in Larger-Scale Settings
In the base setting (Table 5), teachers have very low train loss and moderate test accuracy. With augmentation (Table 5), train loss increases while test accuracy improves, as expected.
Having established how augmentation changes the teacher regime, we now examine the students under the same settings (teacher seed 0). In the base case, KD with 0.9 improves over SIDDO by at most (Activation Distance), (Rank Disagreement), (Prediction Disagreement), and (JS Divergence) (Table 6) – statistically significant (Appendix Table 29) but marginal in magnitude. Under augmentation, KD with 0.9 improves by (Activation Distance), (Rank Disagreement), (Prediction Disagreement), and (JS Divergence) (Table 7). In both base and augmented settings, the best test loss/accuracy occurs under random control distillation, indicating that improved performance does not require a meaningful teacher signal.
@@
Figure 11(a) shows the corresponding prediction agreement deltas (KD vs. best control). At , students trained from augmented teachers increase incorrect agreement from (base) to , far outpacing the increase in correct agreement. Thus, inducing higher teacher train loss via augmentation reliably amplifies asymmetric incorrect transfer, consistent with the SVHN findings and our regularisation view of KD with the novel insight of negative asymmetric transfer.
4.3 Function Transfer Across Modalities
@@
We test the generality of our findings beyond images by evaluating KD on audio (UrbanSound8K, SpeechCommands) and language (Tiny Shakespeare). Across modalities, the same pattern holds: when transfer is non-marginal (per functional similarity metrics), it is asymmetric: students preferentially increase agreement with the teacher on incorrect predictions, and this imbalance strengthens as the teacher weight increases. Below we show the VGGish architecture on the audio datasets and the NanoGPT on Tiny Shakespeare.
In Figure 11(b), a clear pattern emerges: when there is considerable knowledge transfer, as evidenced by results across functional similarity metrics (Appendix Sections G and H), an asymmetric relationship becomes evident in the nature of the transfer. Specifically, student models receive significantly more transfer of the teacher model’s incorrect predictions than its correct predictions, with this imbalance scaling linearly as the weighting on the teacher outputs increases. These results highlight the generality of our understanding of knowledge distillation as a data-dependent regulariser with a negative asymmetric payoff.
4.4 Adversarial Transfer (Language): Targeted error copying
To move beyond aggregate functional similarity, we test whether KD copies a specific erroneous behaviour from its teacher. Informed by the Zipf’s Law distribution (Piantadosi, 2014) of the Tiny Shakespeare dataset as seen in Figure 11, we construct an adversarially biased Tiny Shakespeare teacher by editing its training corpus so that every instance of “the” is replaced with “tha”, a sequence that does not occur in the clean dataset (Appendix H.2, Table 100). This induces a stable bias to complete “th_” as “tha” rather than “the”, while the teacher’s overall performance on clean data remains comparable to standard models (Table 101). We then distil this teacher at and compare against our two controls (SIDDO and RCD) under identical training conditions.
| Predicted Word | Teacher | SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 |
| tha | 454 | 105.90 4.168 | 106.00 3.046 | 199.10 13.391 | 436.20 7.984 | 104.60 3.898 | 114.80 3.056 | 126.90 8.068 |
| the | 285 | 665.10 7.675 | 675.50 10.228 | 583.40 17.536 | 343.60 6.358 | 668.80 12.713 | 712.50 12.480 | 826.30 20.203 |
On clean evaluation prompts containing “th_”, we measure how often models complete to “tha” versus “the” and aggregate results per teacher seed, as seen for teacher seed 0 in Table 8 (with seeds 1-2 in Appendix Tables 103 and 104). KD, particularly at higher , markedly increases the rate of “tha” completions and suppresses “the” relative to both controls, demonstrating that KD can selectively copy a targeted error pattern even when overall behaviour appears benign. This experiment adds causal evidence that KD transmits specific erroneous structure, not merely broad functional alignment, sharpening the safety implication of our main findings: practitioners may unknowingly inherit unintended behaviours from the teacher, reinforcing our characterisation of KD as a data-dependent regulariser with a negative asymmetric payoff. Full details and per-seed statistics are provided in Appendix H.2.
4.5 Distillation Scaling Laws
The preceding sections established when KD transfers knowledge, this transfer is negatively asymmetric. We now ask how these effects evolve with capacity. Distillation Scaling Laws (DSL) (Busbridge et al., 2025) quantify how much student loss changes with compute, teacher quality, and model size. Our study complements DSL by asking how much is transferred as capacity grows: we decompose the distillation signal into correct vs. incorrect teacher–student agreement, offering a mechanistic reading of the “teacher quality” term and explaining negative-transfer regimes that are invisible from loss alone. Concretely, on Tiny Shakespeare we sweep student width from to in steps under a fixed-epoch budget matched to the teacher, using the same optimiser. For each width and , we measure the change in correct and incorrect agreement relative to the best control baseline (means SEM over 10 runs; teacher seed 0). In Figure 12 three core trends emerge which are descirbed below.
1) Student capacity helps, but mainly by amplifying the teacher’s mistakes: as width increases, both correct and incorrect agreement rise, yet the incorrect column grows much faster (from to width at , correct agreement vs. incorrect ).
2) Small students suffer negative transfer: at - width, the incorrect boost is comparable to or larger than the correct.
3) Increasing capacity unlocks more of the distillation signal: however what flows first, and most strongly, is the teacher’s error pattern.
Taken together, these scaling results reveal what is driving the loss curves: KD acts as a data-dependent regulariser with a negative asymmetric payoff, and scaling up the student amplifies the asymmetry of transfer.
5 Gradient-Level Explanation of Asymmetric Transfer
We now provide a concise theoretical explanation for the observed asymmetric error transfer in KD, and in Appendix B extend our functional analysis with information-theoretic and geometric perspectives to quantify when and how alignment with the teacher becomes harmful. These analyses clarify the risks of distillation, especially in safety-critical settings.
Consider the standard KD objective:
where and are the student and teacher logits, respectively. The per-logit gradient is:
with and .
When is the correct class (), the includes both supervision and teacher alignment. But when is an incorrect class (), the gradient reduces to:
This pulls the student toward any non-zero mass the teacher places on that incorrect class. The strength of this pull scales with and the teacher’s own loss. This simple derivation explains our central finding: when the teacher is imperfect, KD disproportionately transfers its errors to the student. The resulting alignment is asymmetric, favouring incorrect predictions. By contrast, if the teacher logits are replaced with a uniform distribution – as in label smoothing (Appendix D) or our random control distillation – the gradient on incorrect classes becomes flat, removing this error-amplifying signal. Empirically, these baselines match or exceed KD’s accuracy, while showing no rise in incorrect agreement. Overall we argue that the observed asymmetric transfer in KD is not incidental but rather emerges directly from the structure of the KD objective.
6 Conclusion
Across controlled self-distillation, small/large-scale settings, cross-modality (image, audio, language), a targeted error test, capacity scaling, standard KD setting with smaller students (Appendix E), and multiple KD variants (Appendix C), KD seldom delivers robust “knowledge transfer”. When transfer occurs, it is typically marginal and inconsistent, and increases with teacher imperfection, amplifying the teacher’s errors more than its correct behaviour (negative asymmetry). By contrast, Random Control Distillation often yields the best loss/accuracy, indicating that reported gains can arise from generic regularisation rather than faithful knowledge transmission. The targeted language experiment confirms KD can copy specific erroneous patterns, and scaling law experiments show capacity amplifies incorrect agreement faster than correct. We therefore reframe KD as a data-dependent regulariser with asymmetric knowledge sharing, with clear safety implications: audit teacher error structure and report functional transfer analyses (correct vs. incorrect agreement) alongside accuracy/loss.
Appendix A Safety Implications of Knowledge Distillation
The insights from our results can be summarised into three key points. 1) knowledge distillation enables statistically significant functional transfer. 2) The accuracy and loss benefits provided by knowledge distillation are often matched or even exceeded by random controls. 3) Knowledge distillation disproportionately transfers incorrect information, with this asymmetry increasing as the proportion of knowledge transfer grows. Considering these findings – particularly points 2 and 3 – Knowledge Distillation raises significant safety concerns. While it is often assumed that knowledge distillation benefits student models, our results challenge this notion by demonstrating a high likelihood that backdoors or harmful artifacts within teacher models could be transferred to student models. We present a concrete case of adversarial transfer facilitated by Knowledge Distillation in Appendix Section H.2. Moreover, we argue that knowledge distillation is not a safe or reliable method. At best, it results in minimal positive transfer, and at worst, it facilitates substantial negative transfer from teacher to student, undermining its practical utility.
Appendix B Extended Functional Analysis: Information-Theoretic and Geometric Perspectives
We apply two additional metrics: Variation of Information (VoI), an information-theoretic measure over discrete labellings that penalises confident mispredictions (Meilă, 2003), and Orthogonal Procrustes Distance (OPD), a geometric alignment metric over output representations (Schönemann, 1966; Ding et al., 2021). We compute VoI and OPD for two representative setups: ResNet18 on SVHN and ResNet50 on TinyImageNet (teacher seed 0). OPD closely tracks trends observed in Activation Distance and JS Divergence, showing decreasing student–teacher discrepancy as increases. VoI generally follows this trend, but diverges in specific cases (high on SVHN) where it increases despite stronger functional alignment. This is not contradictory: VoI penalises confident yet incorrect predictions more heavily than other metrics. Its rise coincides with the strongest observed increase in student–teacher agreement on incorrect predictions, providing further evidence of KD’s asymmetric payoff. Overall, OPD confirms alignment, but VoI reveals when that alignment corresponds to the transfer of incorrect information. Moreover, this behaviour is predicted by our gradient-based analysis: the per-logit gradient under KD pulls the student toward the teacher’s incorrect predictions with strength proportional to and to the teacher’s own loss. VoI captures the cost of absorbing these errors, providing an explicit signal of negative information transfer. OPD, meanwhile, confirms that overall alignment is occurring, but not necessarily to the student’s benefit.
@@ Appendix B Extended Functional Analysis: Information-Theoretic and Geometric Perspectives
| Activation Distance | 0.063 +- 0.002 | 0.064 +- 0.001 | 0.060 +- 0.001 | 0.059 +- 0.001 | 0.144 +- 0.001 | 0.493 +- 0.000 | 0.849 +- 0.000 |
| Rank Disagreement | 0.696 +- 0.003 | 0.688 +- 0.004 | 0.684 +- 0.003 | 0.681 +- 0.003 | 0.800 +- 0.002 | 0.798 +- 0.002 | 0.802 +- 0.003 |
| Prediction Disagreement | 0.045 +- 0.001 | 0.046 +- 0.001 | 0.043 +- 0.001 | 0.042 +- 0.001 | 0.042 +- 0.001 | 0.043 +- 0.001 | 0.046 +- 0.001 |
| JS Divergence | 0.025 +- 0.001 | 0.025 +- 0.001 | 0.023 +- 0.001 | 0.022 +- 0.000 | 0.053 +- 0.000 | 0.201 +- 0.000 | 0.431 +- 0.000 |
| Information Variation | 0.550 +- 0.051 | 0.588 +- 0.049 | 0.594 +- 0.024 | 0.614 +- 0.018 | 0.638 +- 0.000 | 0.638 +- 0.000 | 0.638 +- 0.000 |
| Procrustes Distance | 0.165 +- 0.003 | 0.168 +- 0.004 | 0.164 +- 0.003 | 0.162 +- 0.005 | 0.291 +- 0.001 | 0.304 +- 0.001 | 0.311 +- 0.003 |
| Accuracy | 0.952 +- 0.001 | 0.951 +- 0.001 | 0.954 +- 0.001 | 0.954 +- 0.001 | 0.957 +- 0.001 | 0.957 +- 0.001 | 0.955 +- 0.001 |
| Loss | 0.385 +- 0.011 | 0.344 +- 0.008 | 0.310 +- 0.006 | 0.293 +- 0.004 | 0.236 +- 0.003 | 0.692 +- 0.001 | 1.698 +- 0.001 |

### v2

@@ 3.3 Hypothesis Testing
To evaluate whether KD increases teacher–student functional similarity beyond controls, we compare KD to SIDDO and RCD on each metric and condition. We use paired tests by matching runs across conditions: for each architecture–dataset–teacher-seed setting and student seed, models share the same initialisation and the same data order, isolating the effect of the distillation signal from stochastic factors. For a given architecture–dataset–teacher-seed setting and KD coefficient , our primary hypotheses against controls are:
KD students, on average, are not more functionally similar to the teacher than control models.
@@
For each metric, we compare the 10 KD runs for each value to the corresponding control values offered by both SIDDO and RCD using one-sided Wilcoxon signed-rank tests (Woolson, 2007). Our scientific question is directional: functional similarity corresponds to improved values in the preferred direction for each metric (lower for distance/disagreement metrics; higher for agreement/accuracy; lower for loss).
To control the expected false discovery rate across the functional similarity metrics within each setting, we apply Benjamini–Hochberg False Discovery Rate (FDR) correction (Yekutieli and Benjamini, 1999) for each in {0.1,0.5 and 0.9} and teacher seed, and report significance after FDR. For accuracy and loss, we keep the significance threshold at . We classify transfer as statistically supported only when i) the corrected significance threshold is satisfied, and ii) the effect is in the expected direction for a majority of the functional similarity metrics under both control comparisons. The control conditions are SIDDO and RCD, and variable of interest is standard KD. We use marginal statistically supported transfer to describe cases where transfer is statistically supported, but the raw transfer metrics remain close to control. This indicates limited shared function, with only weak student–teacher agreement on both correct and incorrect teacher predictions.
@@ 4 Results and Discussion
We examine functional transfer facilitated by knowledge distillation across three modalities: computer vision, audio, and language, presenting results in this order. In Section 4.1 we study self-distillation, where the student has the capacity to match the teacher’s function, and in the main body we analyse teacher seed 0. Concretely, we report SVHN (ResNet18) and TinyImageNet (ResNet50) for vision, SpeechCommands and UrbanSound8K (VGGish) for audio, and TinyShakespeare (Nano-GPT/Pico-GPT) for language. Full results for remaining teacher seeds (1–2), additional datasets, and architectures are provided in the appendix (SVHN: Appendix D.2; TinyImageNet: Appendix D.1; audio: Appendix E; language: Appendix F). We choose widely used benchmarks across vision, audio, and language to test the generality of our findings across modalities.
Following self-distillation, Section 4.2 studies standard distillation with smaller students and examines how negative asymmetric transfer changes with student capacity, including a scaling-law analysis (Busbridge et al., 2025) (additional seeds in Appendix C). We then test KD variants via feature-map matching (Section 4.3), and probe how transfer depends on KD hyperparameters by studying the impact of temperature (Section 4.4). Our final experiments illustrate downstream implications of negative asymmetric transfer via a controlled targeted error-transfer setting, showing that KD can copy a specific erroneous pattern from an adversarially modified teacher (Section 4.5). Section 5 summarises our main empirical findings with a gradient-level mechanistic explanation, showing that negative asymmetric behaviour arises naturally from the standard KD objective under teacher error mass and motivating safety-relevant auditing of teacher error modes. In Appendix A, we show that these conclusions are consistent under information-theoretic and geometric measures in addition to our functional similarity metrics, and Appendix B shows that RCD is equivalent to label smoothing while preserving the KD optimisation shown in equation (1).
Training details for all settings are also provided in the appendix. Unless specified otherwise, we report means and 1 SEM over 10 runs per teacher seed and condition in tables and figures. We use the terms statistically supported and marginal as defined in Section 3.3.
4.1 Self-Distillation
We begin with the computer vision domain (Section 4.1.1). In small-scale settings, we observe that when functional transfer is non-marginal and statistically supported, transferred agreement is consistently asymmetric toward the teacher’s errors. We then validate these findings at larger scale on TinyImageNet. In the base setting, transfer is marginal; however, increasing teacher generalisation performance (via augmentation) amplifies both functional transfer and the negative asymmetry of transfer. Finally, we evaluate audio and language (Sections 4.1.2 and 4.1.3), where we observe stronger functional transfer coupled with increased negative asymmetry.
4.1.1 Computer Vision
For the small-scale experiments, we use SVHN with ResNet18. Here, we observe that KD yields statistically supported functional similarity at high values, but the magnitude and asymmetry of transfer vary across teacher seeds. When transfer is non-marginal, we observe a systematic increase in student–teacher agreement on incorrect predictions relative to correct ones. Table 2 shows teacher variability: train losses of , , and with a generalisation gap of for seeds 0, 1, and 2, respectively.
@@ 4.1.1 Computer Vision
When analysing the functional transfer for teacher seed 0 (Table 3), we observe that KD across values yields only marginal transfer relative to the SIDDO and RCD baselines. Notably, the best test loss and accuracy (Table 3) are achieved by random control distillation, reducing confidence that KD’s performance gains arise from meaningful knowledge transfer and instead supporting the view of KD as a data-dependent regulariser. For the highest-train-loss teacher (seed 2), KD produces statistically supported functional transfer across metrics at most values (Table 4), with the exception of Prediction Disagreement at . This transfer coincides with a large asymmetric payoff in prediction agreement toward the teacher’s incorrect predictions (Figure 3). The lowest-train-loss teacher (seed 1) shows no statistically supported transfer at and only partial transfer at (again, excluding Prediction Disagreement). For seed 0 (intermediate train loss), we observe statistically supported transfer at and , accompanied by asymmetric incorrect agreement (Figure 3).
@@
| KD 0.5 | ✗✗✓ | ✗✗✓ | ✗✗✓ | ✗✗✓ | ✗✗✗ | ✗✗✓ |
| KD 0.9 | ✓✗✓ | ✓✗✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✓ |
Across seeds, higher teacher train loss is associated with stronger, more asymmetric, functional transfer, consistent with a teacher that deviates more from ground-truth labels, thereby exposing students to incorrect structure that is preferentially transferred under KD.
Function Transfer in Larger-Scale Settings
To understand how these results scale, we next study TinyImageNet with ResNet50. In the base setting, KD produces statistically supported but marginal functional gains relative to SIDDO in Table 6; the corresponding prediction agreement shows no clear preference toward correct or incorrect agreement in Figure 4. Motivated by the SVHN analysis, we increase the teacher train loss via data augmentation using RandAugment (Cubuk et al., 2020) with the default settings, and examine the consequences for functional transfer and asymmetry. Here, we increase generalisation performance while increasing the teacher’s loss (Table 5) to study the nature of knowledge transfer at increased dataset and architecture complexity.
@@
In the base setting (Table 5), teachers have very low train loss and moderate test accuracy; with augmentation, train loss increases while test accuracy improves, as expected.
@@
| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✓✗ | ✗✗✗ |
| KD 0.5 | ✓✗✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✓✗ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
Having established how augmentation changes the teacher regime, we now examine how this impacts functional similarity between students and teachers (teacher seed 0). In the base case (Table 6), KD with improves functional similarity over SIDDO by at most (Activation Distance), (Rank Disagreement), (Prediction Disagreement), and (JS Divergence). Table 7 shows that this qualifies as statistically supported transfer; however, it is marginal in magnitude. Under augmentation (Table 8), we observe substantially larger functional similarity gains: KD with improves by (Activation Distance), (Rank Disagreement), (Prediction Disagreement), and (JS Divergence), corroborating the hypothesis that increased teacher error mass is associated with stronger functional transfer. Relative to the base teachers, augmentation increases train loss while improving test accuracy, and we observe larger functional alignment under KD in this regime. Notably, in both the base and augmented settings, the best test loss/accuracy occurs under random control distillation, indicating that improved performance does not require a meaningful teacher signal.
@@
Figure 4 shows the corresponding prediction agreement deltas (KD vs. best control), highlighting exactly what agreement is transferred between the teacher and the students in the base and augmented condition. At , students trained from augmented teachers increase incorrect agreement from (base) to , far outpacing the increase in correct agreement. Thus, inducing higher teacher train loss via augmentation reliably amplifies asymmetric incorrect transfer, consistent with the SVHN findings and our regularisation view of KD, with the additional, novel insight of negative asymmetric transfer.
4.1.2 Audio
@@
We test our findings beyond computer vision by evaluating KD on two audio datasets: UrbanSound8K and SpeechCommands. The trend is consistent: when knowledge transfer is non-marginal (according to functional similarity metrics), it is asymmetric. Students preferentially increase agreement with the teacher on incorrect predictions, and this negative asymmetric transfer intensifies as the teacher weight increases. Below, we show results of the VGGish architecture on both audio datasets.
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| JS Divergence () | 0.085 | 0.063 | 0.060 | 0.059 | 0.120 | 0.274 | 0.512 |
| Accuracy () | 0.870 | 0.886 | 0.887 | 0.884 | 0.892 | 0.882 | 0.844 |
| Loss () | 1.076 | 0.669 | 0.564 | 0.553 | 0.565 | 1.103 | 2.366 |
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✓ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ |
For the VGGish on SpeechCommands, Table 10 shows substantially higher knowledge transfer compared to the controls than those observed in SVHN and base case for TinyImageNet. Notably, we observe an approximate reduction in divergence across all similarity metrics, alongside a monotonic increase in functional similarity as increases. Table 11 confirms that for all values and teacher seeds, knowledge transfer is statistically supported for KD. The type and strength of this transfer, shown in Figure 5, reveals that the increased knowledge transfer in SpeechCommands corresponds to significant negative asymmetric transfer, further strengthening the link between statistically supported knowledge transfer via improved functional similarity and severe negative asymmetric transfer. Given, that RCD at = 0.1 provides the best performance, in this instance it is clear that functional alignment and transfer is not responsible for performance benefits.
@@
| Accuracy () | 0.795 | 0.787 | 0.796 | 0.796 | 0.808 | 0.806 | 0.585 |
| Loss () | 2.813 | 2.460 | 2.225 | 2.089 | 0.730 | 1.085 | 2.059 |
| KD 0.1 | ✗✗✓ | ✗✗✓ | ✗✗✓ | ✗✗✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✓✓ | ✓✓✓ | ✗✓✓ | ✓✗✓ | ✗✗✗ | ✗✗✗ |
For VGGish on UrbanSound8K, we observe less knowledge transfer relative to the baselines in Table 12 than on SpeechCommands in Table 10. Once again, RCD 0.1 provides the best overall performance, but here RCD 0.9 causes a large reduction in accuracy on this dataset. When analysing the prediction agreement distribution in Figure 6, it can still be observed that negative asymmetric transfer remains the most prominent form of knowledge transfer. In this figure, we also present the prediction agreement results without RCD 0.9, since its very low accuracy biases the results toward underestimating the degree of asymmetric transfer from KD. When RCD 0.9 is removed, the degree of negative asymmetric transfer becomes more consistent with the statistically supported functional similarity we record across all similarity metrics and seeds. As a result, the findings in audio confirm that negative asymmetric transfer is prominent across modalities and that functional transfer facilitated by KD does not provide the best performance and transfers incorrect teacher representations at a disproportionate rate.
4.1.3 Language
The final modality we consider is language. Having observed negative asymmetric transfer in computer vision and audio, we test whether the same phenomenon occurs in language models. In this section, we study the Nano-GPT architecture on TinyShakespeare, which is where we observe the largest statistically supported knowledge transfer across similarity metrics, coupled with the strongest negative asymmetric transfer.